# Universal Vessel Segmentation for Multi-Modality Retinal Images

### v2

@@ Abstract
We identify two major limitations in the existing studies on retinal vessel segmentation: (1) Most existing works are restricted to one modality, i.e, the Color Fundus (CF). However, multi-modality retinal images are used every day in the study of retina and retinal diseases, and the study of vessel segmentation on the other modalities is scarce; (2) Even though a small amount of works extended their experiments to limited new modalities such as the Multi-Color Scanning Laser Ophthalmoscopy (MC), these works still require finetuning a separate model for the new modality. The finetuning will require extra training data, which is difficult to acquire. In this work, we present a foundational universal vessel segmentation model (UVSM) for multi-modality retinal images. Not only do we perform the study on a much wider range of modalities, but we also propose a universal model to segment the vessels in all these commonly-used modalities. Despite being much more versatile comparing with existing methods, our universal model still demonstrates comparable performance with the state-of-the-art finetuned methods. To the best of our knowledge, this is the first work that achieves cross-modality retinal vessel segmentation and also the first work to study retinal vessel segmentation in some novel modalities.  [^1] Our codes and model will be available at the time of publication. We are also actively looking for approaches to release the three new vessel segmentation datasets. However, since retinal images and retinal vessels contain sensitive personal information, we can not make promise at this moment.
@@
retina, retinal images, vessel segmentation, image translation, topology-preserving image segmentation, foundation model.
@@ I Background and Introduction
Retinal vessel segmentation is a fundamental task in retinal image processing. It has multiple useful applications, including retinal image registration [1, 2], human biometric identification [3], retinal artery-vein classification [4] and tree topology analysis in digital images [5]. Clinically, the evaluation of retinal vessels is also important since the vessel features are predictive in several systemic diseases [6, 7, 8] and vessel segmentation is the key towards the automation of the evaluation. However, most works in retinal vessel segmentation focus on only the Color Fundus (CF) images. CF uses white visible light to illuminate the retina and the reflected light is used for generating the image. Nevertheless, although CF was the most widely-used modality in the past, it requires dilation before the imaging and is uncomfortable for the patients. Additionally, the pathological information provided by CF is limited. Therefore, in more recent years, CF is used less frequently and emerging new imaging techniques become increasingly popular. These modalities include (Fig. 1): (1) Multi-Color Scanning Laser Ophthalmoscopy (MC), which uses more intense laser beam and allow exposing structures in deeper layers of the retina; (2) Florescence Angiography (FA), which uses florescence dye to illuminate the retinal vessels and are commonly used in the diagnosis of diseases related to neovascularization; (3) Fundus AutoFlorescence (FAF), which does not rely on external light source but uses natural fluorescence of the retina and provides insights into the health and metabolism of the retinal pigment epithelium (RPE) and photoreceptor cells; (4) Infrared/Near-Infrared Reflectance (IR/NIR), which uses Infrared or Near-Infrared lights to illuminate the image and allows seeing deeper features on the retina such as RPE abnormalities. (5) Optical Coherence Tomography Angiography (OCTA), which uses low-coherence interferometry to measure the backscattered light from retinal layers. It is primarily used to image the micro-vascular structure on the retina. Nevertheless, OCTA has a much smaller imaging area than the other modalities and has significantly different retinal vessel morphology and topology. Therefore, in this work we focus on the first four modalities and the CF.
@@
There are limited previous works which are extended beyond CF modality [11, 12, 13]. In addition to lacking comprehensive modalities (compared to our approach), one still needs to finetune a separate model on the target modality where each model only works for one image modality. We instead aim to build a single universal model that achieves robust segmentation for all the commonly-used modalities. The model will serve as an important tool in cross-modality retinal image study, which becomes increasingly important in modern research of the retina. To achieve the universal segmentation with limited labeled data, we develop our model in two major steps (Fig. 2). Firstly, we propose to use image translation as a preliminary domain adaptation method and train a model that translate input images from arbitrary modalities and cameras into a uniform Topcon (The Topcon TRC-50DX series is explicitly referred to as the gold standard in color fundus photography) CF image modality. We accessed abundant of de-identified multi-modality retinal images from multiple widely-used cameras (Topcon, Heidelberg Spectralis, Optos) at the Jacobs Retina Center at Shiley Eye Institute, University of California, San Diego (UCSD). Those images allow us to train a useful image translation model in a self-supervised approach. Furthermore, we adopt data augmentation to account for the slight domain difference with the other CF cameras which our center does not have, including Canon and Zeiss. Secondly, we train a segmentation model on the Topcon CF domain to perform the segmentation. However, image translation can not fully eliminate the domain gaps between different modalities given such a difficult many-to-one translation task. To solve this problem, we propose to use a topology-aware segmentation method that learns topological features of the vessels in addition to their conventional pixel-level features. The topology-aware segmentation methods were initially proposed to improve the topological accuracy of the segmentation of tubular structures within a certain type of image. We extend these methods to our task and found that they improve the topological accuracy in the segmentation (e.g., preserving vessel continuity, branching and loop features), as well as they force the segmentation to be topologically correct so when images with slight domain gap are inputted, the model can still make the correct segmentation.
@@
Our universal vessel segmentation model is evaluated on 7 diverse datasets (including 3 new datasets we prepared for this work) of the 5 most commonly-used retinal image modalities (CF, FA, FAF, MC, IR/NIR) taken from 5 most commonly-used cameras (Topcon, Canon, Zeiss, Heidelberg Spectralis, Optos). Our universal model is compared with state-of-the-art methods which are finetuned on each of the 7 datasets, respectively, and is shown to achieve comparable performance. Extensive ablation studies were also performed to investigate how different design choices in the proposed pipeline affects the performance of the universal vessel segmentation. The main contributions of this paper are:
  • We propose a foundational universal vessel segmentation model (UVSM) for multi-modality retinal images which can perform robust vessel segmentation for all the commonly-used modalities and cameras using one single model without providing any modality information.
@@
  • We propose to use a topology-aware image segmentation method to train our segmentation model to allow more robust segmentation across modalities. Furthermore, we discover an additional merit of the topology-aware segmentation loss functions in segmenting images with domain gaps in addition to their conventional applications.
@@
  • We provide a refined version of the widely-used DRIVE dataset with better topological accuracy in the vessel labels (to be discussed in section IV-A2).
@@ II-A Retinal Vessel Segmentation
Retinal vessel segmentation classifies each pixel in the retinal images into foreground (vessel) and background (non-vessel). Current works focus on segmenting vessels in the CF images, where various machine learning-based or non-machine learning-based methods were proposed. A majority of the early methods are based on the matched filter [14, 15, 16], which convolve retinal images with predefined or learned kernels and then adaptively threshold the image to obtain the vessel segmentation map. Another type of methods uses tracking algorithms to mathematically estimate the growth of the retinal vessel trees from seed points [17, 18]. The segmentation can then be derived by extending around the estimated vessel centerlines using the estimated vessel diameters. Some methods also adopt morphological image processing methods such as the top-hat operation to extract the vessel or as a post-processing step [19, 20]. Furthermore, numerous traditional machine learning-based methods were also applied in retinal vessel segmentation such as the support vector machine [21], random forest [22] and Adaboost [23] for supervised methods, and Gaussian mixture model [24], Fuzzy C-Means [25] for unsupervised methods. In more recent years, deep learning-based methods dominate the literature. They mainly rely on the Convolutional Neural Networks (CNN) as a backbone model due to the existence of vast fine details and the spread-out characteristic of the vessels. An introductory work is the Deep Retinal Image Understanding (DRIU) [26], which adopts a VGG [27] network structure as encoder and assigns specialized output layers to the embedded features to segment the vessels and the optic disc from the retina. Based on a widely-used U-Net [28] network structure and to seek further improvement for retinal vessel segmentation, [29, 30, 31, 32, 33] proposed to add various global reasoning modules at the bottom of the network. [32, 33, 34, 35, 31] proposed to add feature aggregation modules in the residual connection at each level of the network. [36, 37] proposed to use deformable convolution where the convolution kernel shapes are learned to adapt to the elongated morphology of the retinal vessels. With similar motivation, [38] proposed an orientation-selective convolution to learn selective receptive fields of the kernel to adapt to the shape of the vessels. [31] proposed to use improved feature fusion methods to better preserve the fine details of vessels in the segmentation. In addition to a single U-Net-like segmentation network, several works [39, 38, 40] proposed to use cascaded networks or a single looped network to iteratively refine the segmentation.
@@ II-C Topology-Aware Image Segmentation
Another type of methods uses strict topological features of the segmentation which are based on the computational topology theory and received more research attentions in recent years. The topological features are mainly based on the persistent homology theory [63, 64, 65, 66, 37, 67, 68, 69]. A revolutionary work that first introduce this concept in image segmenation is [63], which computes the persistent features from the ground truth and prediction, matching the features using a Wasserstein distance [70] and using the distance as the loss function. The following works share similar idea. [64] also uses the Wasserstein distance but apply the method on 3D point cloud data. [67] proposed to match the longest persistent barcodes in the prediction to the ground truth. [37] proposed to use a Hausdorff distance [71] to match the persistent diagrams to improve the sensitivity of the matching to the outliers. [69] and [68] both aim to compute more precise persistent feature matching, as in earlier methods the actual matching results are highly inaccurate due to only considering the global topological features. [69] and [68] instead allow the loss to act more selectively on the topological features thus improve the efficiency of the topological loss function. [68] proposed to use an Induced Matching [72] method which constructs a common ambient space between the ground truth and prediction and maps the persistent features in between to find a better matching. However, the computational cost of this method is very high, especially for large images, where typical retinal images usually have high resolutions. Alternatively, [69] proposed to leverage the spatial locations of the persistent features referencing their critical cell locations for more faithful persistent feature matching and is computationally much more efficient than [68].
@@ III-A Overview
Following the pre-processing, we train an domain adaptation model for multi-modality retinal images. Although several domain adaptation methods were proposed for retinal vessel segmentation, such as using the pseudo-label methods [39], adversarial learning [33] and asymmetrical maximum classifier discrepancy [73], they were proposed only to address the domain gap between different datasets (from different cameras) in CF images. Moreover, these methods still require unsupervised finetuning on the target domain images, which deviate from our goal of universal segmentation. Therefore, in this work we propose to use image translation as the domain adaptation method, which allows us to train a single model for all the modalities.
@@
After the image translation model is trained, we use it to translate our segmentation training set (19 images from the DRIVE [10] training set with vessel labels) to the Topcon CF domain. The translated images are combined with the original images as an augmented training set to train the downstream segmentation network. Nevertheless, since we are training a difficult many-to-one (FA, FAF, MC, IR/NIR to CF) translation, the translation model cannot perfectly address the domain gap between the non-CF images and the CF images (Fig. 12). We therefore propose to further solve this problem by introducing a topology-aware segmentation method. In addition to a conventional pixel-wise loss function (BCELoss), the segmentation network is jointly trained with two topological loss functions.
@@ III-B Image Translation Network
Multiple GAN-based or Diffusion model-based image translation methods are explored for our task. For Diffusion-based method, we experiment with StyleDiffusion [58] and Unpaired Neuron Schrodinger Bridge (UNSB) [53]. For GAN-based method, we experiment with CycleGAN [41]. We will show in section IV-D2 that CycleGAN, despite being a classic method, outperforms the other two most recent methods in most datasets. Therefore, we pick CycleGAN as the backbone of our image translation model. Consequently, the following sections discuss the implementation based on CycleGAN and we refer the readers to [58, 53] for details about the other two methods and to section IV-D2 for our implementation details.
@@
To train the translation network, we treat all the non-CF images (FA, FAF, MC, IR/NIR) as image type A and the Topcon CF images as image type B. At each iteration, a pair of image A batch and image B batch is drawn from the training set. As illustrated by Fig. 3, generator A is trained to translate all the type A images (non-CF) to a fake type B (CF) and generator B is used to translate type B to a fake type A, which is a neutral non-CF domain. Both generator A and B are learned with adversarial learning [74, 75], where the generator and discriminator are alternatively updated. At each iteration, given the image batch $I_{A}$, the generators (A+B) are first updated with the objectives (discriminators frozen):
@@
$\textit{L}_{adv.-G_{A}}=\left|D_{A}(G_{A}(I_{A}))-1\right|^{2}$ (2)
@@
$\textit{L}_{cycle-A}=\left|I_{A}-G_{B}(G_{A}(I_{A}))\right|$ (3)
@@
$\textit{L}_{identity-A}=\left|I_{A}-G_{B}(I_{A})\right|$ (4)
@@
$\textit{L}_{adv.-D_{A}}=\frac{1}{2}\left|D_{A}(I_{B})-1\right|^{2}+\frac{1}{2}\left|D_{A}(G_{A}(I_{A}))\right|^{2}$ (5)

### v3

@@ Abstract
We identify two major limitations in the existing studies on retinal vessel segmentation: (1) Most existing works are restricted to one modality, i.e, the Color Fundus (CF). However, multi-modality retinal images are used every day in the study of retina and diagnosis of retinal diseases, and the study of vessel segmentation on the other modalities is scarce; (2) Even though a few works extended their experiments to limited new modalities such as the Multi-Color Scanning Laser Ophthalmoscopy (MC), these works still require finetuning a separate model for the new modality. The finetuning will require extra training data, which is difficult to acquire. In this work, we present a novel universal vessel segmentation model (UVSM) for multi-modality retinal images. Not only do we perform the study on a much wider range of modalities, but we also propose a universal model to segment the vessels in all these commonly-used modalities. Despite being much more versatile comparing with existing methods, our universal model still demonstrates comparable performance with the state-of-the-art finetuned methods. To the best of our knowledge, this is the first work that achieves modality-agnostic retinal vessel segmentation and also the first work that studies retinal vessel segmentation in some novel modalities.  [^1] Our codes and model will be available at the time of publication. We are also actively looking for approaches to release the three new vessel segmentation datasets. However, since retinal images and retinal vessels contain sensitive personal information, we can not make promise at this moment.
@@
retina, retinal image processing, vessel segmentation, image translation, topological learning.
@@ I Background and Introduction
Retinal vessel segmentation is a fundamental task in retinal image processing. It has multiple useful applications, including retinal image registration [1, 2], human biometric identification [3], retinal artery-vein classification [4] and tree topology analysis in digital images [5]. Clinically, the evaluation of retinal vessels is also important since the vessel features are predictive in several systemic diseases [6, 7, 8] and vessel segmentation is the key towards the automation of the evaluation. However, most works in retinal vessel segmentation focus on only the Color Fundus (CF) images. CF uses white visible light to illuminate the retina and the reflected light is used for generating the image. Nevertheless, although CF was the most widely-used modality in the past, it requires dilation before imaging and is uncomfortable for the patients. Additionally, the pathological information provided by CF is limited. Therefore, in more recent years, CF is used less frequently and emerging new imaging techniques become increasingly popular. These modalities include (Fig. 1): (1) Multi-Color Scanning Laser Ophthalmoscopy (MC), which uses more intense laser beam and allow exposing structures in deeper layers of the retina; (2) Florescence Angiography (FA), which uses florescence dye to illuminate the retinal vessels and are commonly used in the diagnosis of diseases related to neovascularization; (3) Fundus AutoFlorescence (FAF), which does not rely on external light source but uses natural fluorescence of the retina and provides insights into the health and metabolism of the retinal pigment epithelium (RPE) and photoreceptor cells; (4) Infrared/Near-Infrared Reflectance (IR/NIR), which uses Infrared or Near-Infrared lights to illuminate the image and allows seeing deeper features on the retina such as RPE abnormalities. (5) Optical Coherence Tomography Angiography (OCTA), which uses low-coherence interferometry to measure the backscattered light from retinal layers. OCTA has a much smaller imaging area than the other modalities and has considerably different retinal vessel morphology and topology.
@@
There are limited previous works which are extended beyond CF modality [11, 12, 13]. In addition to lacking comprehensive modalities (compared to our approach), one still needs to finetune a separate model on the target modality where each model only works for one image modality. We instead aim to build a single universal model that achieves robust segmentation for all the commonly-used modalities. The model could serve as an important tool in cross-modality retinal image study, which becomes increasingly important in modern research of the retina. To achieve the universal segmentation with limited labeled data, we develop our model in two major steps (Fig. 2). Firstly, we propose to use image translation as a preliminary domain adaptation method and train a model that translate input images from arbitrary modalities and cameras into a uniform Topcon (The Topcon TRC-50DX series is explicitly referred to as the gold standard in color fundus photography) CF image modality. We accessed abundant of de-identified multi-modality retinal images from multiple widely-used cameras (Topcon, Heidelberg Spectralis, Optos) at the Jacobs Retina Center at Shiley Eye Institute, University of California, San Diego (UCSD). Those images allow us to train a useful image translation model in a self-supervised approach. Furthermore, we adopt data augmentation to account for the slight domain difference with the other CF cameras which our center does not have, including Canon and Zeiss. Secondly, we train a segmentation model on the Topcon CF domain to perform the segmentation. However, image translation cannot fully eliminate the domain gaps between different modalities given such difficult many-to-one translation task. To solve this problem, we propose to use a topology-aware segmentation method that learns topological features of the vessels in addition to their conventional pixel-level features. The topology-aware segmentation methods were initially proposed to improve the topological accuracy of the segmentation of curvilinear structures. We go beyond their conventional applications and propose the topological learning as a novel domain adaptation method for retinal vessel segmentation, in a way that forces the segmentation to be topologically correct so when images from different domains are inputted, the model can still output robust segmentation.
@@
Our universal vessel segmentation model is evaluated on 10 diverse datasets (including 3 new datasets we prepared for this work) of the 6 most commonly-used retinal image modalities (CF, FA, FAF, MC, IR/NIR, OCTA) taken from 6 most commonly-used camera systems (Topcon, Canon, Zeiss, Heidelberg Spectralis, Optos, Optovue). Our universal model is compared with state-of-the-art methods which are finetuned on each of the 7 datasets, respectively, and is shown to achieve comparable performance. Extensive ablation studies were also performed to investigate how different components in the proposed pipeline affects the performance of the universal vessel segmentation. The main contributions of this paper are:
  • We propose the universal retinal vessel segmentation task. To the best of our knowledge, this is the first work that treat retinal vessel segmentation as a modality-agnostic segmentation task covering multiple imaging modalities and camera systems, reflecting the real-world need for generalizable retinal vessel segmentation solutions.
  • We propose the first universal (retinal) vessel segmentation model (UVSM) for multi-modality retinal images based on image translation and image topological learning.
@@
  • We propose topological feature learning as a novel domain adaptation strategy for multi-modal retinal vessel segmentation. We show that by learning the topological features of the vessels, the segmentation model becomes much more robust to inputs from different domains (modalities and imaging systems).
@@
  • We provide a refined version of the widely-used DRIVE dataset with improved topological accuracy in the annotation (to be discussed in section IV-A2).
@@ II-A Retinal Vessel Segmentation
Retinal vessel segmentation is a specialized task that general-purpose foundational segmentation models [14, 15, 16] struggle to handle effectively. Current works in retinal vessel segmentation focus on segmenting vessels in Color Fundus retinal images, where various machine learning-based or non-machine learning-based methods were proposed. A majority of the early methods are based on the matched filter [17, 18, 19], which convolve retinal images with predefined or learned kernels and then adaptively threshold the image to obtain the vessel segmentation map. Another type of methods uses tracking algorithms to mathematically estimate the growth of the retinal vessel trees from seed points [20, 21]. The segmentation can then be derived by extending around the estimated vessel centerlines using the estimated vessel diameters. Some methods also adopt morphological image processing methods such as the top-hat operation to extract the vessel or as a post-processing step [22, 23]. Furthermore, numerous traditional machine learning-based methods were also applied in retinal vessel segmentation such as the support vector machine [24], random forest [25] and Adaboost [26] for supervised methods, and Gaussian mixture model [27], Fuzzy C-Means [28] for unsupervised methods. In more recent years, deep learning-based methods dominate the literature. They mainly rely on the Convolutional Neural Networks (CNN) as a backbone model due to the existence of vast fine details and the spread-out characteristic of the vessels. An introductory work is the Deep Retinal Image Understanding (DRIU) [29], which adopts a VGG [30] network structure as encoder and assigns specialized output layers to the embedded features to segment the vessels and the optic disc from the retina. Based on a widely-used U-Net [31] network structure and to seek further improvement for retinal vessel segmentation, [32, 33, 34, 35, 36] proposed to add various global reasoning modules at the bottom of the network. [35, 36, 37, 38, 34] proposed to add feature aggregation modules in the residual connection at each level of the network. [39, 40] proposed to use deformable convolution where the convolution kernel shapes are learned to adapt to the elongated morphology of the retinal vessels. With similar motivation, [41] proposed an orientation-selective convolution to learn selective receptive fields of the kernel to adapt to the shape of the vessels. [34] proposed to use improved feature fusion methods to better preserve the fine details of vessels in the segmentation. In addition to a single U-Net-like segmentation network, several works [42, 41, 43] proposed to use cascaded networks or a single looped network to iteratively refine the segmentation.
@@ II-C Topology-Aware Image Segmentation
Another type of methods uses strict topological features of the segmentation which are based on the computational topology theory and received more research attentions in recent years. The topological features are mainly based on the persistent homology theory [66, 67, 68, 69, 40, 70, 71, 72]. A revolutionary work that first introduces this concept in image segmenation is [66], which computes the persistent features from the ground truth and prediction, matching the features using a Wasserstein distance [73] and using the distance as the loss function. The following works share similar idea. [67] also uses the Wasserstein distance but apply the method on 3D point cloud data. [70] proposed to match the longest persistent barcodes in the prediction to the ground truth. [40] proposed to use a Hausdorff distance [74] to match the persistent diagrams to improve the sensitivity of the matching to the outliers. [72] and [71] both aim to compute more precise persistent feature matching, as in earlier methods the actual matching results are highly inaccurate due to only considering the global topological features. [72] and [71] instead allow the loss to act more selectively on the topological features thus improve the efficiency of the topological loss function. [71] proposed to use an Induced Matching [75] method which constructs a common ambient space between the ground truth and prediction and maps the persistent features in between to find a better matching. However, the computational cost of this method is very high, especially for large images, where typical retinal images usually have high resolutions. Alternatively, [72] proposed to leverage the spatial locations of the persistent features referencing their critical cell locations for more faithful persistent feature matching and is computationally much more efficient than [71].
@@ III-A Overview
Following pre-processing, we train a domain adaptation model for multi-modality retinal images. Although several domain adaptation methods were proposed for retinal vessel segmentation, such as using the pseudo-label methods [42, 76], teacher-student network [77], adversarial learning [36, 76] and asymmetrical maximum classifier discrepancy [78], they were proposed only to address the domain gap between different datasets (from different cameras) in CF images. Moreover, these methods still require unsupervised finetuning on the target domain images, which deviate from our goal of universal segmentation. Therefore, in this work we propose to use image translation as the preliminary domain adaptation method, which allows us to train a single model for all the modalities.
@@
After the image translation model is trained, we use it to translate our segmentation training set (19 images from the DRIVE [10] training set with vessel labels) to the Topcon CF domain. The translated images are combined with the original images as an augmented training set to train the downstream segmentation network. Nevertheless, since we are training a difficult many-to-one (FA, FAF, MC, IR/NIR to CF) translation, the translation model cannot perfectly address the domain gap between the non-CF images and the CF images (Fig. 13). To solve this problem, we propose topology-aware segmentation as a novel domain adaptation method. We show that by learning additional topological features of the vessels, the segmentation model becomes much more robust to inputs from different domains. In practice, in addition to a conventional pixel-wise loss function (BCELoss), the segmentation network is jointly trained with two topological loss functions.
@@ III-B Image Translation Network
Multiple GAN-based or Diffusion model-based image translation methods are explored for our task. For Diffusion-based method, we experiment with StyleDiffusion [61] and Unpaired Neuron Schrodinger Bridge (UNSB) [56]. For GAN-based method, we experiment with CycleGAN [44]. We will show in section IV-E2 that CycleGAN, despite being a classic method, outperforms the other two most recent methods in most datasets. Therefore, we pick CycleGAN as the backbone of our image translation model. Consequently, the following sections discuss the implementation based on CycleGAN and we refer the readers to [61, 56] for details about the other two methods and to section IV-E2 for our implementation details.
@@
To train the translation network, we treat all the non-CF images (FA, FAF, MC, IR/NIR) as image type A and the Topcon CF images as image type B. At each iteration, a pair of image A batch and image B batch is selected from the training set. As illustrated by Fig. 3, generator A is trained to translate all the type A images (non-CF) to a fake type B (CF) and generator B is used to translate type B to a fake type A, which is a neutral non-CF domain. Both generator A and B are learned with adversarial learning [79, 80], where the generator and discriminator are alternatively updated. At each iteration, given the image batch $I_{A}$, the generators (A+B) are first updated with the objectives (discriminators frozen):
@@
$\textit{L}_{adv.-G_{A}}=\left|D_{A}(G_{A}(I_{A}))-1\right|^{2}_{2}$ (2)
@@
$\textit{L}_{cycle-A}=\left|I_{A}-G_{B}(G_{A}(I_{A}))\right|_{2}$ (3)
@@
$\textit{L}_{identity-A}=\left|I_{A}-G_{B}(I_{A})\right|_{2}$ (4)
@@
$\textit{L}_{adv.-D_{A}}=\frac{1}{2}\left|D_{A}(I_{B})-1\right|^{2}_{2}+\frac{1}{2}\left|D_{A}(G_{A}(I_{A}))\right|^{2}_{2}$ (5)